# Branching Flows: Discrete, Continuous, and Manifold Flow Matching with Splits and Deletions

### v1

@@ A.1.1 CTMC Equivalence with Counting Event Times
Conditioned on $X_{t}=x_{t}$, there are exactly $n-x_{t}$ remaining jumps. Denote $R:=n-x_{t}$. It holds that
@@
Let $T_{t}^{+}:=\min\{T_{i}:T_{i}>t\}$ denote the arrival time of the next jump at time $t$. Condition on the specific at risk index set $S\subset\{1,\ldots,n\}$ with $|S|=R$. Then for any $i\in S$,
$p_{\delta}:=\mathbb{P}(t<T_{i}\leq t+\delta\mid T_{i}>t)=1-\frac{S(t+\delta)}{S(t)}=1-\frac{S(t)+\delta S^{\prime}(t)+o(\delta)}{S(t)}=h(t)\delta+o(\delta),$
@@
$Q_{t}^{z_{c}}(x_{t}+1|x_{t})$ $=(x_{1}-x_{t})h(t)I_{\{x_{t}<z_{c}\}},$
and $Q_{t}^{z_{c}}(x_{t}|x_{t})=-Q_{t}(x_{t}+1|x_{t}),$ with $Q_{t}^{z_{c}}(\cdot|x_{t})=0$ otherwise. In other words, if the event times are distributed according to $H$, then counting the events that have occurred up until the current time out of $z_{c}$ total events corresponds to the counting flow process conditioned on $z_{c}$, starting at zero.
A.1.2 Interarrival Times
Theorem 3 (Interarrival time distribution).
@@ A.1.2 Interarrival Times
where $F_{H}^{-1}(y):=\inf\{y:F_{H}(y)\geq(u)\}$ is the quantile function for $H$. In particular, if $U\sim\mathrm{Unif}(0,1)$, then
@@
We find the quantile function $F_{V_{t}}^{-1}(u)=\inf\{y:F_{V_{t}}(y)\geq u\}$. It holds that
@@
$W_{t}\overset{d}{=}F_{H}^{-1}\left(1-S_{H}(t)(1-U)^{\frac{1}{z_{c}-x_{t}}}\right)$
@@ A.1.3 Marginal Generator
and $Q_{t}^{z_{c}}(x_{t}|x_{t})=-Q_{t}(x_{t}+1|x_{t})$, with $Q_{t}^{z_{c}}(\cdot|x_{t})=0$ otherwise — where $h(t)$ is the hazard of a hazard distribution $H$ supported on $[0,1]$. We will recast this in terms of a time-dependent jump kernel (see Appendix D.1). This specifies a conditional process $X_{t}|Z_{c}=z_{c}$.
@@
corresponding to jumping from $x_{t}$ to $x_{t}+1$ with rate $h(t)(z_{c}-x_{t})$. Using the notion of a time-dependent linear parametrization (BilleraTimedep), we can parametrize the infinitesimal generator corresponding to the jump process above by $R_{t}^{z_{c}}(x)=z_{c}-x$, and have a valid conditional generator matching loss be given by
@@
where $R_{t}^{\theta}(x_{t})$ is emitted by the model for each realisation $x_{t}$ of $X_{t}$, and $D$ is a Bregman divergence. Note that the loss is minimized by the parameterization of the marginal generator, which coincides by linearity with the posterior averaged parameterization of the conditional generator
@@
A.2 Deletion Process
For the deletion process, we simply note that both the waiting time law and the marginal rates are a special case of the Counting process when the counts are either 0 (present) or 1 (deleted), and an increment is interpreted as a deletion.
As justified in BilleraTimedep, whenever the hazard rate multiplier can be interpreted as a probability, i.e. is in the range $[0,1]$, a valid conditional generator matching loss is found in the cross-entropy loss:
$L_{\mathrm{cgm}}(\theta)$ $=-\mathbb{E}_{t\sim\mathrm{\mathcal{D}[0,1]},Z\sim p_{Z},X_{t}\sim p_{t|Z}(dx|z)}[\rho_{t}^{Z}(X_{t})\log\rho_{t}^{\theta}(X_{t})+(1-\rho_{t}^{Z}(X_{t}))\log(1-\rho_{t}^{\theta}(X_{t}))]$
where $\rho_{t}^{Z}$ corresponds to the ground truth probability of a deletion, in this case either zero or one, and $\rho_{t}^{\theta}$ corresponds to the predicted probability of a deletion.
Appendix B Conditional and Marginal Paths for a Noisy Discrete Process
Here, we use an extension of Discrete Flow Matching (DFM Gat2024DiscreteFM) to allow time intervals $t_{0},t_{1}]$.
Let $F_{1}$ and $F_{2}$ be CDFs of some continuous distributions supported on $[0,1]$, and let $\omega_{u}\in[0,1)$. Then we define the schedulers:
With the schedulers above, we consider the interpolant path
$p_{t}(x|x_{t_{0}},x_{t_{1}})=\kappa_{1}(t)\delta_{x_{t_{1}}}(x)+\kappa_{2}(t)p_{u}(x)+\kappa_{3}(t)\delta_{x_{t_{0}}}(x)$
for a uniform weighting term $p_{u}$. This choice enforces $\ell_{\kappa}:=\mathrm{argmin}_{j\in[m]}\frac{\dot{\kappa}_{t}^{j}}{\kappa_{t}^{j}}=3$. In the following, we denote $\kappa_{t}^{j}:=\kappa_{j}(t).$ Define the interpolant path
which follows from applying DFM Theorem 3 on the interval $t_{0},1]$, In particular, let $\alpha_{1}=\kappa_{t}^{1}-\kappa_{t_{0}}^{1}\frac{\kappa_{t}^{3}}{\kappa_{t_{0}}^{3}}$, $\alpha_{2}=\kappa_{t}^{2}-\kappa_{t_{0}}^{2}\frac{\kappa_{t}^{3}}{\kappa_{t_{0}}^{3}}$ and $\alpha_{3}=\frac{\kappa_{t}^{3}}{\kappa^{3}_{t_{0}}}$. It can be seen that $\ell_{\alpha}:=\mathrm{argmin}_{j\in[m]}\frac{\dot{\alpha}_{t}^{j}}{\alpha_{t}^{j}}=3$, along with $\alpha_{1},\alpha_{2},\alpha_{3}\geq 0$ and $\sum_{i=1}^{3}\alpha_{i}=1$. The expression for the rates thus follows from
$=\dot{\kappa}_{t}^{1}-\kappa_{t}^{1}\frac{\dot{\kappa}^{3}_{t}}{\kappa^{3}_{t}}$
and analagously
Appendix C The Ornstein-Uhlenbeck process with Time Dependent Diffusion Coefficient
@@
where $v_{t}$ is the time-dependent infinitesimal variance (so the diffusion coefficient is $\sqrt{v_{t}}$). We will, for some $x_{1}$, consider the time-inhomogeneous OU-bridge $X_{t}\mid X_{1}=x_{1}$ where the unconditional process $X_{t}$ solves
@@
Appendix D Branching Flows in Generator Matching
D.1 The Infinitesimal Generator of Jump Processes
@@
which is a probability measure on $S\setminus\{x\}$. Now we see how specifying a jump kernel $Q_{t}(dy;x)$ describes a jump process. With the hazard/rate of $\lambda_{t}(x)$, we will jump to a point $y\sim J_{t}(dy;x)$. The sampling procedure is thus as follows:
$(X_{t+\Delta t}|X_{t}=x_{t})=\begin{cases}X_{t}&\text{with probability }1-\Delta t\lambda_{t}(x_{t})+o(\Delta t)\\ \sim J_{t}(dy|x_{t})&\text{with probability }\Delta t\lambda_{t}(x_{t})+o(\Delta t)\end{cases}.$
@@
As in Linear Parametrization and Loss (LPAL) (BilleraTimedep) we consider atomic time-dependent jump kernels:
@@
D.2 The Branching Infinitesimal Generator
We begin by noting that more complex augmentation schemes than the one introduced in Section 2.2 are possible. In general, one can transport from $\tilde{p}_{0}$ to $q:=\Phi_{\sharp}\tilde{p}_{1}$ by performing generator matching to transport from $\tilde{p}_{0}$ to $\tilde{p}_{1}$, and then pushing samples from $\tilde{p}_{1}$ through $\Phi$ to obtain samples from $q$. In this section, we simply denote $\mathcal{E}$ for a general component-wise state space, augmented or not.
@@ D.2 The Branching Infinitesimal Generator
Let $S:=\bigsqcup_{n\geq 1}\mathcal{E}^{n}=\{(x,n):x\in n\geq 1,\mathcal{E}^{n},\}$. For this section and elsewhere in the paper, we often denote elements $(x,n)\in S$ simply by $x\in S$ for convenience. We let $h_{\mathrm{inc}}(t)>0$ and $h_{\mathrm{del}}(t)>0$ be overall hazard multipliers such that $h_{\mathrm{inc}}(t)\to\infty$ and $h_{\mathrm{del}}(t)\to\infty$ as $t\nearrow 1$, corresponding to an overall rate multiplier for split increments and an overall rate multiplier for deletions, as in Section 2.3.1 and Section 2.3.2. Suppose that splitting of the $i$’th component out of $n$-components occurs with rate $R_{t,i}(x,n)h_{\mathrm{inc}}(t)$ and that deletion of the $i$’th component out of $n$-components occurs with rate $\rho_{t,i}(x,n)h_{\mathrm{del}}(t)$. Here, $R_{t,i}(x,n)\geq 0$ is intended to indicate the ‘remaining number of splits of the $i$’th component’ and $\rho_{t,i}(x,n)\in[0,1]$ is intended to indicate the ‘probability of deleting the $i$’th component’.
In this section, we arrive at a time- and state varying linear parametrization for the infinitesimal generator of our branching process. The notion of a linear parametrization of a generator, introduced in GM and further discussed in FMG, was clarified to be able to vary with both time and state in LPAL. We find a time- and state varying linear parametrization of the branching generator of the form
$\mathcal{L}^{\mathrm{branch}}_{t}f(x,n)$ $=\langle\mathcal{K}_{t,n}^{\mathrm{split}}f(x),R_{t}(x)\rangle_{t,n}+\langle\mathcal{K}_{t,n}^{\mathrm{del}}f(x),\rho_{t}(x)\rangle_{t,n}+\langle\mathcal{K}_{t,n}^{\mathrm{base}}f(x),F_{t}^{\mathrm{base}}(x)\rangle_{t,(x,n)}^{\mathrm{base}}.$
@@
Using an atomic time-dependent jump kernel as in LPAL, splits and deletions as above correspond to
$Q_{t}(dy|(x,n))$ $=\sum_{i=1}^{n}h(t)R_{t,i}(x)\delta_{\mathrm{split}_{i}(x)}(dy)+\sum_{i=1}^{n}h(t)\rho_{t,i}(x)\delta_{\mathrm{del}_{i}(x)}(dy)$
$=h(t)\cdot\left(\sum_{i=1}^{n}R_{t,i}(x)\delta_{\mathrm{split}_{i}(x)}(dy)+\sum_{i=1}^{n}\rho_{t,i}(x)\delta_{\mathrm{del}_{i}(x)}(dy)\right).$
For each $f\in\mathcal{T}(S)$, viewing $f(x,n)=f(x)$ where it is convenient, the infinitesimal generator for jumps becomes
$\mathcal{L}_{t}^{\mathrm{jump}}f(x,n)$ $=\int_{E}(f(y)-f(x))Q_{t,(x,n)}(dy|(x,n))$
$=\int_{E}\textstyle\bigg((f(y)-f(x))\left[h_{\mathrm{inc}}(t)\cdot\sum_{i=1}^{n}R_{t,i}(x)\delta_{\mathrm{split}_{i}(x)}(dy)+h_{\mathrm{del}}(t)\cdot\sum_{i=1}^{n}\rho_{t,i}(x)\delta_{\mathrm{del}_{i}(x)}(dy)\right]\bigg)$
$=h_{\mathrm{inc}}(t)\cdot\sum_{i=1}^{n}\Big((f(\mathrm{split}_{i}(x))-f(x))R_{t,i}(x)\Big)+h_{\mathrm{del}}(t)\cdot\sum_{i=1}^{n}\Big(f(\mathrm{del}_{i}(x))-f(x))\rho_{t,i}(x)\Big).$
We express the jump generator in terms of a time-and-state varying linear parametrization of the generator, as is justified in LPAL, one for splitting and one for deletion. For $v,w\in\mathbb{R}^{n}$, we define the time-varying inner products
$\langle v,w\rangle_{t,n}^{\mathrm{split}}{}:=h_{\mathrm{inc}}(t)\langle v,w\rangle_{\mathbb{R}^{n}}\qquad\text{and}\qquad\langle v,w\rangle_{t,n}^{\mathrm{del}}:=h_{\mathrm{del}}(t)\langle v,w\rangle_{\mathbb{R}^{n}}.$
Note that those define bona fide inner products, since $h_{\mathrm{inc}}(t)>0$ and $h_{\mathrm{del}}(t)>0$. This allows us to express
$\mathcal{L}_{t}^{\mathrm{jump}}f(x,n)=\langle\mathcal{K}_{t,n}^{\mathrm{split}}f(x),R_{t}(x,n)\rangle_{t,n}^{\mathrm{split}}+\langle\mathcal{K}_{t,n}^{\mathrm{del}}f(x),\rho_{t}(x,n)\rangle_{t,n}.$
In the above, we set:
$R_{t}(x,n)$ $=(R_{t,i}(x))_{i=1}^{n}$
@@
$\mathcal{L}_{t}^{\mathrm{base}}f(x,n)=\langle\mathcal{K}_{t,n}^{\mathrm{base}}f(x),F_{t}^{\mathrm{base}}(x,n)\rangle_{t,(x,n)}^{\mathrm{base}}$

### v3

@@ B.1.1 Conditional distributions
Conditioned on $X_{t}=x_{t}$, there are exactly $n-x_{t}$ remaining events. Denote $R:=n-x_{t}$. It holds that
@@
Let $T_{t}^{+}:=\min\{T_{i}:T_{i}>t\}$ denote the arrival time of the next jump at time $t$. Condition on the specific at risk index set $A=\{i:1\leq i\leq n,\ T_{i}>t\}$, corresponding to $|A|=R$ remaining events. Then for any $i\in A$ we have
$p_{\delta}:=\mathbb{P}(t<T_{i}\leq t+\delta\mid T_{i}>t)$ $=1-\frac{S(t+\delta)}{S(t)}$
$=1-\frac{S(t)+\delta S^{\prime}(t)+o(\delta)}{S(t)}=h(t)\delta+o(\delta),$
@@
$Q_{t}^{z_{c}}(x_{t}+1|x_{t})$ $=(z_{c}-x_{t})h(t)I_{\{x_{t}<z_{c}\}},$
and $Q_{t}^{z_{c}}(x_{t}|x_{t})=-Q_{t}^{z_{c}}(x_{t}+1|x_{t}),$ with $Q_{t}^{z_{c}}(\cdot|x_{t})=0$ otherwise. In other words, if the event times are distributed according to $H$, then counting the events that have occurred up until the current time out of $z_{c}$ total events corresponds to the counting flow process conditioned on $z_{c}$, starting at zero.
B.1.2 Interarrival Times
Theorem B.1.3.
@@ B.1.2 Interarrival Times
where $F_{H}^{-1}(u):=\inf\{y:F_{H}(y)\geq u\}$ is the quantile function for $H$. In particular, if $U\sim\mathrm{Unif}(0,1)$, then
@@
The quantile function is $F_{V_{t}}^{-1}(u)=\inf\{y:F_{V_{t}}(y)\geq u\}$. It holds that
@@
$W_{t}\overset{d}{=}F_{H}^{-1}\left(1-S_{H}(t)(1-U)^{\frac{1}{z_{c}-x_{t}}}\right)-t$
@@ B.1.3 Marginal Generator
and $Q_{t}^{z_{c}}(x_{t}|x_{t})=-Q_{t}^{z_{c}}(x_{t}+1|x_{t})$, with $Q_{t}^{z_{c}}(\cdot|x_{t})=0$ otherwise. We will recast this in terms of a time-dependent jump kernel (see Appendix E.1). This specifies a conditional process $X_{t}|Z_{c}=z_{c}$.
@@
corresponding to jumping from $x_{t}$ to $x_{t}+1$ with rate $h(t)(z_{c}-x_{t})$. It holds that all remaining events will occur by $t=1$ almost surely. Using the notion of a time-dependent linear parametrization (billera2025timedependentlossreweighting), we can parametrize the infinitesimal generator corresponding to the jump process above by $R_{t}^{z_{c}}(x)=z_{c}-x$, and have a valid conditional generator matching loss be given by
@@
where $R_{t}^{\theta}(x_{t})$ is emitted by the model for each realization $x_{t}$ of $X_{t}$, and $D$ is a Bregman divergence. Note that the loss is minimized by the parameterization of the marginal generator, which coincides by linearity with the posterior averaged parameterization of the conditional generator
@@
B.2 Deletion Flows
We also describe a CTMC in the state space $S=\{0,1\}$, where 0 is ‘present’ and 1 is ‘deleted’. Let $H_{\mathrm{del}}$ be a hazard distribution supported on $[0,1]$ with no atom at $1$, with its associated hazard rate
$h_{\mathrm{del}}(t)=\frac{f_{H_{\mathrm{del}}}(t)}{1-F_{H_{\mathrm{del}}}(t)}$
We represent the conditional probability of a deletion by $t=1$ with $Z_{d}\in\{0,1\}$. This determines the conditional rate $Q_{t}^{Z_{d}}(x|x_{t})$, defined by:
$Q^{Z_{d}}_{t}(1|0)=Z_{d}h_{\mathrm{del}}(t)\quad\text{and}\quad Q_{t}^{Z_{d}}(0|0)=-Q^{Z_{d}}_{t}(1|0),$
along with $Q^{Z_{d}}_{t}(\cdot|x_{t})=0$ for all other values of $x_{t}$. Conditional on $Z_{d}=1$, corresponding to a deleted element, the process $X_{t}|(Z_{d}=1)$ will terminate in deletion by $t=1$ with probability $1$. Let $\rho_{t}^{z_{d}}=z_{d}$ and $\rho_{t}^{\theta}$ be a neural network’s prediction of the probability of deletion by time $t$. Then
is a valid, per-term, conditional generator matching loss under an appropriately scaled time-dependent linear parameterization (billera2025timedependentlossreweighting).
Remark B.2.1.
The generator and the interarrival times obtained from having one or zero remaining events for counting flows will coincide for deletion flows.
B.2.1 Branching Flows
We construct Branching Flows with split and deletion events. The hazard rate of split events is defined identically to that of the increments in Counting Flows. For splits, the waiting time until the next event will coincide with that in Theorem B.1.3, where the number of remaining split events ahead of each node corresponds to the number of remaining increments in Counting Flows. Deletions are then exactly as in Deletion Flows.
Appendix C Branch tracking, splits, and deletions
and analogously define $\tilde{x}_{t}^{(i)}\triangleright\downarrow$.
$\operatorname{split}_{i}(\tilde{x}_{t}):=(\tilde{x}_{t}^{(1)},\ldots,(\tilde{x}_{t}^{(i)}\triangleright\uparrow),\ (\tilde{x}_{t}^{(i)}\triangleright\downarrow),\ldots,\tilde{x}_{t}^{(n)})).$
that deletes the element $\tilde{x}^{(i)}$ at index $i$.
Appendix D Base processes
D.1 The Ornstein-Uhlenbeck process with Time Dependent Diffusion Coefficient
@@
where $\theta>0$ and $v_{t}$ is the time-dependent infinitesimal variance (so the diffusion coefficient is $\sqrt{v_{t}}$). We will, for some $x_{1}$, consider the time-inhomogeneous OU-bridge $X_{t}\mid X_{1}=x_{1}$ where the unconditional process $X_{t}$ solves
@@
D.2 Conditional and Marginal Paths for a Noisy Discrete Process
In what follows, we use an extension of Discrete Flow Matching (DFM Gat2024DiscreteFM) to allow time intervals $t_{0},t_{1}]$. Let $F_{1}$ and $F_{2}$ be CDFs of some continuous distributions supported on $[0,1]$, and let $\omega_{u}\in[0,1)$. Then we define the schedulers
It can be shown that our choice of schedulers enforces $\ell_{\kappa}:=\mathrm{argmin}_{j\in[m]}\frac{\dot{\kappa}_{t}^{j}}{\kappa_{t}^{j}}=3$. With these schedulers, define the interpolant path
which follows from applying DFM Theorem 3 on the interval $t_{0},1]$. To see this, define $\alpha_{1}=\kappa_{t}^{1}-\kappa_{t_{0}}^{1}\frac{\kappa_{t}^{3}}{\kappa_{t_{0}}^{3}}$, $\alpha_{2}=\kappa_{t}^{2}-\kappa_{t_{0}}^{2}\frac{\kappa_{t}^{3}}{\kappa_{t_{0}}^{3}}$ and $\alpha_{3}=\frac{\kappa_{t}^{3}}{\kappa^{3}_{t_{0}}}$. It can be seen that $\ell_{\alpha}:=\mathrm{argmin}_{j\in[m]}\frac{\dot{\alpha}_{t}^{j}}{\alpha_{t}^{j}}=3$, along with $\alpha_{1},\alpha_{2},\alpha_{3}\geq 0$ and $\sum_{i=1}^{3}\alpha_{i}=1$. The expression for the rates follows from
$=\dot{\kappa}_{t}^{1}-\kappa_{t}^{1}\frac{\dot{\kappa}^{3}_{t}}{\kappa^{3}_{t}}$
and analogously
Appendix E Branching Flows in Generator Matching
E.1 The Infinitesimal Generator of Jump Processes
@@
which is a probability measure on $S\setminus\{x\}$. Now we see how specifying a jump kernel $Q_{t}(dy;x)$ describes a jump process. With the hazard/rate of $\lambda_{t}(x)$, we will jump to a point $y\sim J_{t}(dy;x)$.
For $\lambda_{t}(x_{t})=0$ the process is constant, and otherwise the sampling procedure is as follows:
$(X_{t+\Delta t}|X_{t}=x_{t})=\begin{cases}X_{t}&\text{with probability }1-\Delta t\lambda_{t}(x_{t})+o(\Delta t),\\ \sim J_{t}(dy;x_{t})&\text{with probability }\Delta t\lambda_{t}(x_{t})+o(\Delta t).\end{cases}$
@@
As in (billera2025timedependentlossreweighting) we consider atomic time-dependent jump kernels:
@@
E.2 The Marginal Branching Infinitesimal Generator Rewrite
@@
Let $S:=\bigsqcup_{n\geq 1}\mathcal{E}^{n}$. We say that $x\in\mathcal{E}^{n}$ has $n$ elements. We let $h_{\mathrm{split}}(t)$ and $h_{\mathrm{del}}(t)$ be overall hazard multipliers as in Appendix B.1, corresponding to an overall rate multiplier for split increments and an overall rate multiplier for deletions. Suppose that splitting of the $i$’th element out of the $n$ elements occurs with rate $h_{\mathrm{split}}(t)\cdot R_{t,i}(x)$ and that deletion of the $i$’th element out of the $n$ elements occurs with rate $h_{\mathrm{del}}(t)\cdot\rho_{t,i}(x)$. Here, $R_{t,i}(x)\geq 0$ is intended to indicate the ‘remaining number of splits of the $i$’th element’ and $\rho_{t,i}(x)\in[0,1]$ is intended to indicate the ‘probability of deleting the $i$’th element’. For $n=1$, we require that the deletion probability is zero so as not to result in an empty sequence. In this section, we arrive at a time- and state varying linear parametrization for the infinitesimal generator of our branching process. The notion of a linear parametrization of a generator, introduced in Holderrieth2024GeneratorMG and further discussed in Lipman2024FlowMG, was clarified to be able to vary with both time and state in billera2025timedependentlossreweighting.
For the subsequent discussion, we consider times $t$ such that $h_{\mathrm{split}}(t)>0$ and $h_{\mathrm{del}}(t)>0$. This is justified because if the hazard rate for split events $h_{\mathrm{split}}(t)$ is zero, there is nothing for the model to learn since in that case it is known a priori that the split rate $h_{\mathrm{split}}(t)\cdot R_{t,i}(x)$ vanishes. Similarly, if the hazard rate for deletion events $h_{\mathrm{del}}(t)$ is zero, the deletion rate $h_{\mathrm{del}}(t)\cdot\rho_{t,i}(x)=0$ is also known a priori.
In what follows, we consider $x\in\mathcal{E}^{n}$ having $n\geq 1$ elements. A time- and state varying linear parametrization of the branching generator is
$\mathcal{L}^{\mathrm{branch}}_{t}f(x)$ $=\langle\mathcal{K}_{t,n}^{\mathrm{split}}f(x),R_{t}(x)\rangle_{t,n}^{\mathrm{split}}+\langle\mathcal{K}_{t,n}^{\mathrm{del}}f(x),\rho_{t}(x)\rangle_{t,n}^{\mathrm{del}}+\langle\mathcal{K}_{t,n}^{\mathrm{base}}f(x),F_{t}^{\mathrm{base}}(x)\rangle_{t,x}^{\mathrm{base}}.$
@@
In the above, using an atomic time-dependent jump kernel as in (billera2025timedependentlossreweighting), splits and deletions correspond to the time-dependent jump kernel
$Q_{t}(dy;x)$ $=\sum_{i=1}^{n}h_{\mathrm{split}}(t)R_{t,i}(x)\delta_{\mathrm{split}_{i}(x)}(dy)+\sum_{i=1}^{n}h_{\mathrm{del}}(t)\rho_{t,i}(x)\delta_{\mathrm{del}_{i}(x)}(dy).$
With this time-dependent jump kernel, the infinitesimal generator for jumps becomes
$\mathcal{L}_{t}^{\mathrm{jump}}f(x)$ $=\int_{S}(f(y)-f(x))Q_{t}(dy;x)$
$=\int_{S}\bigg((f(y)-f(x))\bigg[h_{\mathrm{split}}(t)\cdot\sum_{i=1}^{n}\Big(R_{t,i}(x)\delta_{\mathrm{split}_{i}(x)}(dy)\Big)$
$\qquad+h_{\mathrm{del}}(t)\cdot\sum_{i=1}^{n}\Big(\rho_{t,i}(x)\delta_{\mathrm{del}_{i}(x)}(dy)\Big)\bigg]\bigg)$
$=h_{\mathrm{split}}(t)\cdot\sum_{i=1}^{n}\Big((f(\mathrm{split}_{i}(x))-f(x))R_{t,i}(x)\Big)$
$\qquad+h_{\mathrm{del}}(t)\cdot\sum_{i=1}^{n}\Big((f(\mathrm{del}_{i}(x))-f(x))\,\rho_{t,i}(x)\Big).$
To linearly parametrize this generator, we define the inner products
$\langle v,w\rangle_{t,n}^{\mathrm{split}}{}:=h_{\mathrm{split}}(t)\langle v,w\rangle_{\mathbb{R}^{n}}\qquad\text{and}\qquad\langle v,w\rangle_{t,n}^{\mathrm{del}}:=h_{\mathrm{del}}(t)\langle v,w\rangle_{\mathbb{R}^{n}}.$
This allows us to linearly parametrize the jump generator with
$\mathcal{L}_{t}^{\mathrm{jump}}f(x)=\langle\mathcal{K}_{t,n}^{\mathrm{split}}f(x),R_{t}(x)\rangle_{t,n}^{\mathrm{split}}+\langle\mathcal{K}_{t,n}^{\mathrm{del}}f(x),\rho_{t}(x)\rangle_{t,n}^{\mathrm{del}},$
where in the above, we set
@@
$\mathcal{L}_{t}^{\mathrm{base}}f(x)=\langle\mathcal{K}_{t,n}^{\mathrm{base}}f(x),F_{t}^{\mathrm{base}}(x)\rangle_{t,x}^{\mathrm{base}}$